# InteractPro: A Unified Framework for Motion-Aware Image Composition

### v2

@@ Abstract
1 Introduction
How can a static object, like a book, be realistically integrated into a background image with a sandcastle, capturing the deformation and compression of the sandcastle under the weight of the book? This scenario presents a significant challenge in achieving realistic image composition—the process of blending a foreground object with a background to create a convincing scene. Although the field has advanced rapidly with developments in diffusion models [35, 10], current techniques face several critical limitations. Techniques like Paint-by-Example [44] and ObjectStitch [37] allow for image-guided editing of specific scene regions using a target image as a template; however, they fall short in producing identity-consistent content, particularly for categories not covered during training. More importantly, although often appearing to be visually pleasant, these methods neglect the interaction between the inserted object and the background scene—an essential aspect for achieving contextual coherence and true realism. This reveals a deeper shortcoming: a lack of motion- and physics-aware modeling. Without capturing the dynamic and structural response between scenes and inserted objects, current methods fall short of delivering truly realistic image compositions. Overall, diffusion-based image models are purely data-driven and often rely on training distributions, limiting their ability to generate physical realism [31, 4, 3, 25], or handle out-of-distribution scenarios for abstract compositions. In addition, they require manual selection of insertion locations, making them less scalable and more labor-intensive for users. Addressing this gap is essential for advancing beyond static visual alignment toward motion-aware, interaction-consistent synthesis.
@@ 1 Introduction
Each module compensates for the other’s limitations: InteractPhys excels at modeling physically grounded interactions—such as deformation or compression—that are difficult to infer visually, but it is limited in highly complex scenes due to the need for explicit 3D representation and simulation. Conversely, InteractMotion handles visually complex or ambiguous scenes more robustly, but often fails to capture implied physical effects—like compression or weight transfer—due to its lack of deep physical grounding [25]. By combining deterministic physical modeling with flexible data-driven synthesis, InteractPro robustly handles both common and out-of-distribution motion-aware composition tasks. To coordinate these strengths, our intelligent planner, InteractPlan, uses GPT-4V [1] with multimodal Chain-of-Thought (CoT) prompting [42, 53, 51], drawing from LLM-assisted frameworks [36, 22, 11]. InteractPlan carefully evaluates the characteristics of the foreground and background images, considering multiple factors such as the type of object interaction and environmental dynamics. By balancing a set of thoughtfully crafted criteria, InteractPlan selects the most appropriate method, ensuring optimal results across a diverse range of image scenarios. InteractPlan also automates optimal placement of foregrounds, ensuring logical and visually consistent integration.
@@
2 Related Works
2.1 Generative Image Composition
Recently, generative image composition techniques [44, 37, 8, 49, 44, 28, 13] have sought to address these multifaceted issues with a comprehensive model, enabling end-to-end generation of image-guided composite images. Prominent among these are Paint-by-Example [44], AnyDoor [8], and ObjectStitch [37], which, despite their progress, struggle with preserving foreground integrity and seamless blending. Crucially, these methods predominantly result in visually pleasant but static compositions, lacking the dynamic interaction and motion coherence. InteractPro stands out by integrating elements seamlessly—even when they come from vastly different visual domains, such as combining photorealistic backgrounds with cartoon elements. It also introduces context-aware motion and structural responses,allowing inserted objects to interact naturally with their environment and enhancing overall realism.
2.2 LLM-Assisted Visual Generations
Large language models (LLMs) like GPT-4 with Vision (GPT-4V) [1] and PaLM [2] excel in various multimodal tasks [23, 39, 26, 16]. A key development in harnessing the potential of LLMs is the introduction of chain-of-thought (CoT) prompting [42, 52], a technique that enhances adaptability to specific tasks. Several approaches has leveraged LLMs as planning component across diverse applications, benefitting from CoT prompting and multimodal reasoning capabilities. [17] shows that LLMs behave like zero-shot planners when they are correctly prompted. [45, 27] showcase LLMs’ planning capabilities by breaking down complex generation tasks into specific subprompts, enabling more precise control over text-to-image and video outputs. [30, 16, 54] utilise the strong planning capabilities of LLMs for Blender script generation and scene understanding, showcasing their potential in creative and technical domains.
2.3 Physics Integrated Representations
In 3D asset creation, physics-based techniques allow the generation of motion driven by physical interactions. [7] achieves joint reconstruction of geometry, appearance, and physical properties for elastic objects through a multiview capture system using compressed air, enabling realistic animations under new physical conditions. [24] and [43] integrate physics simulations with NeRF and 3D Gaussian frameworks, respectively, to produce physically accurate motion. Building on the [43] approach, [50] and [26] further integrate the optimization of material properties into physics-based 3D Gaussians by leveraging pre-trained video generation models. We extend MPM into image composition task and enhance the capabilities of [26] for more physically coherent results.
3 Method
@@ 3.1 Preliminary: Physics3D
3.2 InteractPlan
We leverage the reasoning and multimodal capabilities of GPT-4V as our planner, InteractPlan, which intelligently selects between InteractPhys (Section 3.3) and InteractMotion (Section 3.4) by analyzing object interactions, environmental effects, material properties, and object complexity. This ensures the application of the most suitable method for realistic, seamless compositions across diverse scenarios.
@@ 3.3 InteractPhys
Precise physics integration. Unlike [26], which treats the entire scene of Gaussian objects as a single entity, our enhanced InteractPhys enables per-part parameter control. For a scene/object with segments, each part is assigned its own physical parameters . These variations allow different materials within an object to exhibit distinct responses to external forces, enhancing realism and physical accuracy.
@@ 3.4 InteractMotion
4 Experiment
4.1 Implementation Details
@@
4.2 Quantitative Results
@@
While standard metrics like LPIPS and CLIP scores are adept at measuring static similarity, they lack the sophistication to differentiate between rudimentary cut-and-paste operations and the more nuanced, motion-aware image compositions characteristic of our work. Recognizing the limitations of traditional evaluation methods, we employ a user survey to measure the impact of our motion-aware compositions. This allows us to gauge user perception, which is pivotal in recognizing the subtleties and dynamism our method infuses into the images, providing a more holistic and suitable evaluation framework for our work.
@@
We present quantitative comparisons using a user survey on 52 participants with existing related works: Anydoor [8], ControlCom [49], Cross domain Composition [13] and Paint by Example (PbE) [44]. We provided participants with a standardized Google Forms questionnaire that elaborated on each criterion with definitions and examples to ensure understanding and consistency in responses. We use 30 groups of images, each group contains two inputs (foreground and background images) and five outputs generated by the different methods mentioned. Half of the our results are generated with InteractMotion and half with InteractPhys. Participants are required to select their most preferred method via Multiple Choice Questions, based on object identity consistency, seamless blending, motion aware coherence, and overall harmony considering the above three criteria. We then collated the votes each method received for each criterion to obtain the quantitative results listed in Table 4.2, which demonstrates alignment between subjective user ratings and our objective visual analysis.
4.3 Qualitative Results
@@
We present qualitative results with the baseline works in Fig. 3 and a detailed explanation on our composite results below. Rows (a-b) are performed with InteractMotion and (c-d) are with InteractPhys. In (a), introducing a few bubbles near a thumb demonstrates dynamic interaction, where the bubbles burst upon contact with the thumb, capturing a moment of delicate rupture. In (b), a cup of water is mixed with a yellow-colored liquid. The resulting composite vividly illustrates the water adopting traces of yellow to indicate the mixture. In (c), only the plant’s branches bend upon impact of the falling object, demonstrating precise, localized force control—while the pot remains unchanged due to its higher mass. This level of targeted physical deformation is difficult to achieve with diffusion-based methods, which often fail to convey the impact of heavy objects on specific parts of a scene. In (d), the car deforms under the weight of the heavy log, illustrating realistic bending behavior consistent with the material properties and physical context.
@@
Fig.4 highlights the novel simulation capabilities of InteractPro with engaging and non-traditional scenarios. In panel (a), a visualization captures a car melting under the harsh desert sun, depicting the effects of exaggerated heat. Panel (b) depicts a simulated plant that appears ordinary but exhibits an unusual response: it shrinks upon contact, mimicking the behavior of sensitive plants. In (c), apple transforms into a pile of slim like substance, settling onto the plant and causing it to lean due to the redistributed weight. The base of the pot remains unaffected, anchored by its higher mass. These examples highlight the framework’s potential to apply physical principles creatively, thereby expanding the boundaries of physical interactions in image compositions that cannot be generalized by diffusion-based models.
We show extra qualitative results and explanation of our motion aware composition results for above Fig.1 in Supplementary. We omit qualitative comparisons for atypical scenes, as some baselines do not support controllable interaction or conditioning to reproduce these behaviors.
4.4 Ablation
@@
5 Conclusion
We present InteractPro, a unified framework for motion-aware image composition that combines explicit physics simulation (InteractPhys) and data-driven motion synthesis (InteractMotion) under a planner-guided system. This hybrid approach enables realistic, controllable object-scene interactions across diverse scenarios, addressing key limitations of existing methods and producing visually coherent, physically plausible compositions.
@@ 5 Conclusion
Appendix A Limitations and Future Works
While InteractPro offers a modular framework for physics-aware visual composition, it is not without limitations. InteractPlan currently relies on heuristic decision logic and operates in a single-agent setting, which may result in suboptimal choices in ambiguous or multi-object scenarios. Future work could explore learning-based or multi-agent planners with visual reasoning capabilities to improve robustness and scalability. InteractPhys, while enabling controllable physical interactions, is limited to scenes and objects that can be lifted into 3D, and currently supports only a small set of simulation materials. Extending it to handle partial geometry, learning-based material inference, or hybrid 2D–3D simulation could broaden its applicability. InteractMotion exhibits variability across seeds, with no guarantees of physical plausibility. Future work could incorporate flow-based guidance, motion regularization, or latent alignment constraints to improve consistency and controllability while preserving generative richness.
Appendix B More Preliminary
@@
Appendix C InteractPlan Details
C.1 Full prompt template for method decision
@@
If InteractMotion chosen: Evaluate the optimal split ratio of the background image and optimal insertion region of foreground object: The split ratio Sratio is (x,y); best region R* for insertion is Region Z. The evaluation is based on this process: Section C.2.
@@
Object shape: Ball — geometrically simple and symmetrical; simplifies collision detection and deformation modeling.
@@
Simulation Complexity:
InteractPhys is for localized physical interactions such as collisions, elastic/plastic deformation, and rigid or semi-rigid body contact. Typically used when it is easy to perform simulation.
InteractMotion is preferred for large-scale transformations, complex topology changes, continuous flows, or phenomena involving optical or dynamic surface behavior. Typically used when it is hard to perform simulation.
Material properties:
InteractPhys excels with granular or deformable materials such as sand or jelly.
InteractMotion handles flow-like effects and phenomena such as surface tension. Typically in fluids and gases.
Environmental effects:
InteractMotion is preferred if mechnical forces such as impact or pressure dominate.
InteractPhys is preferred if factors like wind, light dynamics are present.
Object shape and structure:
Simple and uniform shapes can be handled by both modules.
Complex, dynamic shapes favor InteractMotion.
@@
C.2 Object Placement Decision For InteractMotion

### v3

@@ Abstract
Index Terms:
Image Composition, Multimodal Planning, Image Synthesis, Motion Aware Composition, Physics-based Image Generation.
I Introduction
How can a static object, like a book, be realistically integrated into a background image with a sandcastle, capturing the deformation and compression of the sandcastle under the weight of the book? This scenario presents a significant challenge in achieving realistic image composition—the process of blending a foreground object with a background to create a convincing scene. Although the field has advanced rapidly with developments in diffusion models [1, 2], current image composition techniques face several critical limitations. Techniques like Paint-by-Example [3] and ObjectStitch [4] allow for image-guided editing of specific scene regions using a target image as a template; however, they fall short in producing identity-consistent content, particularly for categories not covered during training. More importantly, although often appearing to be visually pleasant, these methods neglect the interaction between the inserted object and the background scene—an essential aspect for achieving contextual coherence and true realism. This reveals a deeper shortcoming: a lack of motion- and physics-aware modeling. Without capturing the dynamic and structural response between scenes and inserted objects, current methods fall short of delivering truly realistic image compositions. Overall, diffusion-based image models are purely data-driven and often rely on training distributions, limiting their ability to generate physical realism [5, 6, 7, 8], or handle out-of-distribution scenarios for abstract compositions. In addition, they require manual selection of insertion locations, making them less scalable and more labor-intensive for users. Addressing this gap is essential for advancing beyond static visual alignment toward motion-aware, interaction-consistent synthesis.
@@ I Introduction
Each module compensates for the other’s limitations: InteractPhys excels at modeling physically grounded interactions—such as deformation or compression—that are difficult to infer visually, but it is limited in highly complex scenes due to the need for explicit 3D representation and simulation. Conversely, InteractMotion handles visually complex or ambiguous scenes more robustly, but often fails to capture implied physical effects—like compression or weight transfer—due to its lack of deep physical grounding [8]. By combining deterministic physical modeling with flexible data-driven synthesis, InteractPro robustly handles both common and out-of-distribution motion-aware composition tasks. To coordinate these strengths, our intelligent planner, InteractPlan, uses GPT-4V [13] with multimodal Chain-of-Thought (CoT) prompting [14, 15, 16], drawing from LLM-assisted frameworks [17, 18, 19, 20, 21]. InteractPlan carefully evaluates the characteristics of the foreground and background images, considering multiple factors such as the type of object interaction and environmental dynamics. By balancing a set of thoughtfully crafted criteria, InteractPlan selects the most appropriate method, ensuring optimal results across a diverse range of image scenarios. InteractPlan also automates optimal placement of foregrounds, ensuring logical and visually consistent integration.
@@
II Related Works
II-A Generative Image Composition
Recently, generative image composition techniques [4, 22, 23, 3, 24, 21, 25, 26] have sought to address these multifaceted issues with a comprehensive model, enabling end-to-end generation of image-guided composite images. Prominent among these are Paint-by-Example [3], AnyDoor [22], and ObjectStitch [4], which, despite their progress, struggle with preserving foreground integrity and seamless blending. Crucially, these methods predominantly result in visually pleasant but static compositions, lacking the dynamic interaction and motion coherence. InteractPro stands out by integrating elements seamlessly—even when they come from vastly different visual domains, such as combining photorealistic backgrounds with cartoon elements. It also introduces context-aware motion and structural responses,allowing inserted objects to interact naturally with their environment and enhancing overall realism.
II-B LLM-Assisted Visual Generations
Large language models (LLMs) like GPT-4 with Vision (GPT-4V) [13] excel in various multimodal tasks [27, 11, 28, 29]. A key development in harnessing the potential of LLMs is the introduction of chain-of-thought (CoT) prompting [14, 30], a technique that enhances adaptability to specific tasks. Several approaches has leveraged LLMs as planning component across diverse applications, benefitting from CoT prompting and multimodal reasoning capabilities. [31] shows that LLMs behave like zero-shot planners when they are correctly prompted. [32, 33] showcase LLMs’ planning capabilities by breaking down complex generation tasks into specific subprompts, enabling more precise control over text-to-image and video outputs. [34, 35] utilise the strong planning capabilities of LLMs for Blender script generation and scene understanding, showcasing their potential in creative and technical domains.
II-C Physics Integrated Representations
In 3D asset creation, physics-based techniques allow the generation of motion driven by physical interactions. [36] achieves joint reconstruction of geometry, appearance, and physical properties for elastic objects through a multiview capture system using compressed air, enabling realistic animations under new physical conditions. [37] and [10] integrate physics simulations with NeRF and 3D Gaussian frameworks, respectively, to produce physically accurate motion. Building on the PhysGaussian [10] approach, [12] and [11] further integrate the optimization of material properties into physics-based 3D Gaussians by leveraging pre-trained video generation models. We extend MPM into image composition task and enhance the capabilities of [11] for more physically coherent results.
III Method
@@ III-A Preliminary: Physics3D
III-B InteractPlan
We leverage the reasoning and multimodal capabilities of GPT-4V as our planner, InteractPlan, which intelligently selects between InteractPhys (Section III-C) and InteractMotion (Section III-D) by analyzing object interactions, environmental effects, material properties, and object complexity. This ensures the application of the most suitable method for realistic, seamless compositions across diverse scenarios.
@@ III-C InteractPhys
Precise physics integration. Unlike Physics3D [11], which treats the entire scene of Gaussian objects as a single entity, our enhanced InteractPhys enables per-part parameter control. For a scene/object with segments, each part is assigned its own physical parameters . These variations allow different materials within an object to exhibit distinct responses to external forces, enhancing realism and physical accuracy.
@@ III-D InteractMotion
IV Experiment
IV-A Implementation Details
@@
IV-B Quantitative Results
| Method | DINO | DINO-fg | DINO-bg | FID | FID-fg | FID-bg |
| Anydoor [22] | 0.9560 | 0.9457 | 0.9751 | 20.7703 | 21.8201 | 13.1277 |
| ControlCom [23] | 0.9318 | 0.9227 | 0.9372 | 32.1804 | 35.2777 | 22.8563 |
| Cross Domain [25] | 0.9613 | 0.9442 | 0.9615 | 22.2624 | 24.8327 | 9.4748 |
| PbE [3] | 0.9399 | 0.9346 | 0.9508 | 24.6288 | 30.2789 | 14.3671 |
| Ours | 0.9726 | 0.9660 | 0.9797 | 20.7736 | 18.4088 | 10.3970 |
| Method | ||||
| Overall | ||||
| Anydoor [22] | 13.46 | 11.54 | 3.85 | 11.54 |
| ControlCom [23] | 7.69 | 3.85 | 7.69 | 7.69 |
| CrossDomain [25] | 13.46 | 11.54 | 11.54 | 9.62 |
| PbE [3] | 3.85 | 5.77 | 3.85 | 5.77 |
| Ours | 61.54 | 67.31 | 73.07 | 65.38 |
For the quantitative evaluations below, we prepare 30 groups of images, each group contains two inputs (foreground and background images) and five outputs generated by the different methods mentioned. Half of the our results are generated with InteractMotion and half with InteractPhys.
We report DINO [54] and Fréchet Inception Distance (FID) [55] scores in Table I, distinguishing between foreground and background consistency. The foreground object should remain semantically identifiable while adapting to its environment, so we report DINO-fg as the primary measure of object identity. The background, by contrast, is expected to remain unchanged, making FID-bg the most suitable metric for distributional fidelity. Our method achieves the best performance on DINO-fg, demonstrating strong preservation of object identity, and ranks second on FID-bg, remaining competitive in background fidelity. For completeness and comparability with prior work, we also report DINO-bg, FID-fg, and overall DINO and FID scores. While these auxiliary metrics are less aligned with our task goals, our method remains consistently among the top two across them.
@@ IV-B Quantitative Results
Importantly, however, while these automatic scores capture background preservation and object identity, they lack the sophistication to differentiate between rudimentary cut-and-paste operations and the more nuanced, motion-aware image compositions characteristic of our work. Recognizing the limitations of traditional evaluation methods, we employ a user survey to measure the impact of our motion-aware compositions. This allows us to gauge user perception, which is pivotal in recognizing the subtleties and dynamism our method infuses into the images, providing a more holistic and suitable evaluation framework for our work.
@@
We present quantitative comparisons using a user survey on 52 participants with existing related works: Anydoor [22], ControlCom [23], Cross domain Composition [25] and Paint by Example (PbE) [3]. We provided participants with a standardized Google Forms questionnaire that elaborated on each criterion with definitions and examples to ensure understanding and consistency in responses. Participants are required to select their most preferred method via Multiple Choice Questions, based on object identity consistency, seamless blending, motion aware coherence, and overall harmony considering the above three criteria. We then collated the votes each method received for each criterion to obtain the quantitative results listed in Table IV-B, which demonstrates alignment between subjective user ratings and our objective visual analysis.
IV-C Qualitative Results
@@
We present qualitative results with the baseline works in Fig. 3 and a detailed explanation on our composite results below. In (a), introducing a few bubbles near a thumb demonstrates dynamic interaction, where the bubbles burst upon contact with the thumb, capturing a moment of delicate rupture. In (b), a cup of water is mixed with a yellow-colored liquid. The resulting composite vividly illustrates the water adopting traces of yellow to indicate the mixture. In (c), only the plant’s branches bend upon impact of the falling object, demonstrating precise, localized force control—while the pot remains unchanged due to its higher mass. This level of targeted physical deformation is difficult to achieve with diffusion-based methods, which often fail to convey the impact of heavy objects on specific parts of a scene. In (d), the car deforms under the weight of the heavy log, illustrating realistic bending behavior consistent with the material properties and physical context.
@@
Fig. 4 highlights the novel simulation capabilities of InteractPro with engaging and non-traditional scenarios. In panel (a), a visualization captures a car melting under the harsh desert sun, depicting the effects of exaggerated heat. Panel (b) depicts a simulated plant that appears ordinary but exhibits an unusual response: it shrinks upon contact, mimicking the behavior of sensitive plants. In (c), apple transforms into a pile of slim like substance, settling onto the plant and causing it to lean due to the redistributed weight. The base of the pot remains unaffected, anchored by its higher mass. These examples highlight the framework’s potential to apply physical principles creatively, thereby expanding the boundaries of physical interactions in image compositions that cannot be generalized by diffusion-based models. Note that we are unable to provide qualitative comparison of baselines against the atypical interaction shown in Fig. 4, as some baselines cannot incorporate additional conditioning—such as text prompts—and can only produce default or typical outputs.
IV-D More explanation of Fig. 1 InteractMotion.
We provide explanation of our motion aware composition InteractMotion in Fig. 1(a-c), with qualitative comparison in Fig. 5. Our method showcases realistic interactions between inserted foreground object and object in given background. In (a), a towel is placed on a rope at the beach, where the image captures the towel swaying in the breeze. Our method inherently incorporates scale and rotation transformations, guiding the towel’s natural motion in response to the wind. Similarly in (b), the flag is inserted into the open air grass field, where wind presence is likely, shows the flag swaying naturally in the air. In (c), placing a glass bottle into the ocean gives seamless integration. InteractMotion effectively positions the bottle within the sea layer, avoiding the disjointed appearance typical of Cut and Paste, or inconsistent identity in other baselines. The realistic distortion from light refraction in water further enhances this effect.
IV-E More explanation of Fig. 1 InteractPhys.
We provide explanation of our physics aware composition InteractPhys in Fig. 1(d-f). In 1(d), the runny egg smoothly spreads over the contours of the hotdog following the shape of hotdog. In 1(e), the branch and leaves of the plant bend under the impact of an apple falling into the plant, while the heavier pot remains stationary. In 1(f), the sandcastle is compressed due to the weight of the book. We also show comparisons with other baselines for these cases in Fig. 6. Other methods are unable to demonstrate the smooth flow of runny egg against the hotdog, or changes the appearance of input egg. For the apple example, others only show fruit hanging on the plant and seems like on different image layer or looking like cut and paste, whereas ours perfectly place the apple in between the branches of the plant, causing the branches to tilt. Other methods are also unable to demonstrate the weight impact of the book, which should cause the sandcastle to be compressed.
IV-F Ablation
@@
V Conclusion
We present InteractPro, a unified framework for motion-aware image composition that bridges physically grounded simulation and data-driven motion synthesis. By combining a planner-guided system with two complementary modules- InteractPhys for explicit physics simulation and InteractMotion for leveraging motion priors from video diffusion models-InteractPro enables realistic and controllable object-scene interactions across broad range of scenarios. Our results show that this hybrid approach effectively addresses limitations in existing methods, producing dynamic compositions that are visually coherent and physically plausible.
Limitations and Future Works. While InteractPro provides a modular framework for physics-aware visual composition, each component has limitations. InteractPlan currently uses heuristic decision rules and operates in a single-agent setting, which can lead to suboptimal choices in ambiguous or multi-object scenarios. Future work could incorporate learning-based or multi-agent planners with visual reasoning to improve robustness and scalability. InteractPhys, though enabling controllable physical interactions, is restricted to scenes and objects that can be reconstructed in 3D and presently supports only a limited set of simulation materials. Extending it to partial geometry, learning-based material inference, or hybrid 2D–3D simulation would broaden its applicability. InteractMotion inherits variability from its underlying video diffusion prior, leading to differences across seeds. Future directions include flow-based guidance, motion regularization, or latent alignment constraints to improve consistency.
@@ V Conclusion
VI More Preliminary
@@
VII InteractPlan Details
VII-A Full prompt template for method decision
@@
If InteractMotion chosen: Evaluate the optimal split ratio of the background image and optimal insertion region of foreground object: The split ratio Sratio is (x,y); best region R* for insertion is Region Z. The evaluation is based on this process: Section VII-B.
@@
Object shape: Ball - geometrically simple and symmetrical; simplifies collision detection and deformation modeling.
@@
Simulation Complexity: InteractPhys is for localized physical interactions such as collisions, elastic/plastic deformation, and rigid or semi-rigid body contact. Typically used when it is easy to perform simulation. InteractMotion is preferred for large-scale transformations, complex topology changes, continuous flows, or phenomena involving optical or dynamic surface behavior. Typically used when it is hard to perform simulation.
Material properties: InteractPhys excels with granular or deformable materials such as sand or jelly. InteractMotion handles flow-like effects and phenomena such as surface tension. Typically in fluids and gases.
Environmental effects: InteractMotion is preferred if mechnical forces such as impact or pressure dominate. InteractPhys is preferred if factors like wind, light dynamics are present.
Object shape and structure: Simple and uniform shapes can be handled by both modules. Complex, dynamic shapes favor InteractMotion.
@@
VII-B Object Placement Decision For InteractMotion